# Formulation of entropy through work by Carnot machine and direct derivation of law of entropy non-decrease from Kelvin principle

### v2

@@ 2.2 Simple systems
irrespective of whether $A_{1}\to A_{2}$ is reversible or irreversible.
Let $A_{0}$ be a reference equilibrium state of the system specified with $(T_{0},V_{0})$. We create the equilibrium state $A_{1}$ from the reference equilibrium state $A_{0}$ reversibly by combining two thermodynamic processes considered in Section 2.1 (Fig. 2). The work extracted from this process is the same as Eq. (11):
$W_{0\to 1}^{\rm rev}=U_{0}-U_{1}-T_{0}(S_{0}-S_{1}).$ (18)
@@
$W_{2\to 0}^{\rm rev}=-W_{0\to 2}^{\rm rev}=-(U_{0}-U_{2}-T_{0}(S_{0}-S_{2})).$ (19)
From Eqs. (17), (18), and (19), we obtain the work extracted from the whole cycle process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\to A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$ as
$W_{\rm cyc}=W_{0\to 1}^{\rm rev}+W_{1\to 2}^{\rm ad}+W_{2\to 0}^{\rm rev}=T_{0}(S_{1}-S_{2}).$ (20)
By applying Kelvin principle Eq. (5) to Eq. (20), we derive the principle of entropy non-decrease Eq. (1).
2.3 Compound systems
We show Eq. (1) from Eq. (5) for a compound system. Here, a compound system is composed of $M$ subsystems in equilibrium, where each subsystem is separated by other subsystems by insulated walls. The state of $i$-th subsystem is specified with $(T_{j},V_{j})$ ($j=1,\cdots,M$). The internal energy, entropy, and volume of the compound system are given by the sum of the ones of subsystems as $U=\sum_{j=1}^{M}U_{j}$, and $S=\sum_{j=1}^{M}S_{j}$, and $V=\sum_{j=1}^{M}V_{j}$.
In the same manner as the simple system, we consider a cycle process that consists of three thermodynamic processes $A_{1}\to A_{2}$, $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}$, and $A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$ as follows. Let $A_{1}$ be the state of a compound system specified with the states of $M_{1}$ subsystems $\{(T_{1,1},V_{1,1}),(T_{1,2},V_{1,2}),\cdots,(T_{1,M_{1}},V_{1,M_{1}})\}$ in equilibrium, where $V_{1}=\sum_{j=1}^{M_{1}}V_{1,j}$ is the volume of the compound system at $A_{1}$. Consider an arbitrary adiabatic process $A_{1}\to A_{2}$, where $A_{2}$ is specified with $M_{2}$ subsystems $\{(T_{2,1},V_{2,1}),(T_{2,2},V_{2,2}),\cdots,(T_{2,M_{2}},V_{2,M_{2}})\}$ in equilibrium. During this process, in addition to the adiabatic operations such as stirring, moving a piston attached to the system, and moving insulated walls partitioning subsystems, the number of subsystems is allowed to change from $M_{1}$ to $M_{2}$ by newly inserting or removing walls partitioning subsystems without any work cost. The work extracted from this process $W_{1\to 2}^{\rm ad}$ is given by
@@ 2.3 Compound systems
Subsequently, we create the state $A_{1}$ from a reference equilibrium state $A_{0}$ of a simple system specified with $(T_{0},V_{0})$. First, we change the volume from $V_{0}$ to $V_{1}$ by a reversible adiabatic process, where we denote the state $(T_{0^{\prime}},V_{1})$ by $A_{0^{\prime}}$. The reversible work extracted from this process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0^{\prime}}$ is given by
@@
Then, we insert insulated walls so that the system is divided into $M_{1}$ subsystems, where each subsystem is in the equilibrium state $A_{0^{\prime},j}$ specified with $(T_{0^{\prime}},V_{1,j})$ ($j=1,\dots,M_{1}$). By operating a Carnot machine between each subsystem and the heat reservoir at temperature $T_{0}$ until it reaches the equilibrium state with $(T_{1,j},V_{1,j})$, the total reversible work extracted from this process $A_{0^{\prime}}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}$ is given by
$W_{0^{\prime}\to 1}^{\rm rev}$ $=\sum_{j=1}^{M_{1}}(U_{0^{\prime},j}-U_{1,j}-T_{0}(S_{0^{\prime},j}-S_{1,j}))$ (23)
$=U_{0^{\prime}}-U_{1}-T_{0}(S_{0^{\prime}}-S_{1}).$
@@
Finally, the reversible work extracted from $A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$ should be calculated. Here, the system at the state $A_{2}$ consists of $M_{2}$ subsystems in equilibrium, whose state we denote by $A_{2,j}$ specified with $(T_{2,j},V_{2,j})$, where $\sum_{j=1}^{M_{2}}V_{j}=V_{2}$. In the similar manner to the calculation of $W_{0\to 1}^{\rm rev}$, we can calculate the reversible work extracted from $A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$ as the minus of $W_{0\to 2}^{\rm rev}$:
$W_{2\to 0}^{\rm rev}=-W_{0\to 2}^{\rm rev}=-(U_{0}-U_{2}-T_{0}(S_{0^{\prime\prime}}-S_{2})),$ (25)
@@
where $S_{0^{\prime\prime}}$ is the entropy at the state $A_{0^{\prime\prime}}$ specified with $(T_{0^{\prime\prime}},V_{2})$ reached from $A_{0}$ in a reversible adiabatic process. Note that $S_{0^{\prime}}=S_{0^{\prime\prime}}$ as $A_{0^{\prime}}$ and $A_{0^{\prime\prime}}$ are on the same adiabatic curve. By summing Eqs. (21), (24), and (25), and using $S_{0^{\prime}}=S_{0^{\prime\prime}}$, we obtain the work extracted from the whole cycle process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\to A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$ as
$W_{\rm cyc}=W_{0\to 1}^{\rm rev}+W_{1\to 2}^{\rm ad}+W_{2\to 0}^{\rm rev}=T_{0}(S_{1}-S_{2}).$ (26)
@@
By applying the Kelvin principle in Eq. (5) to Eq. (26), we derive the principle of entropy non-decrease in Eq. (1) for the compound system.
3 Discussion
3.1 Gyftopoulos and Beretta entropy
@@
It should be noted that our derivation of entropy and its non-decrease in Sec. 2 shares some similarities with the approach by Gyftopoulos and Beretta (GB) [5]. By using notations in our formulation, their entropy is defined through the difference between works extracted from processes with and without the use of a heat reservoir at temperature $T_{0}$ [7]:
$S_{1}=S_{0}+\frac{W_{0\to 1}^{\rm rev}-W_{0\to 1}^{\rm ad}}{T_{0}},\ S_{2}=S_{0}+\frac{W_{0\to 2}^{\rm rev}-W_{0\to 2}^{\rm ad}}{T_{0}}.$ (27)
@@
This definition is equivalent to Eq. (10) and is derived from it with the use of Eq. (6) by noting $W_{0\to 1}^{\rm ad}=U_{0}-U_{1}$, but the one focusing on work rather than heat. This is a natural definition as Eq. (11) suggests that the entropy serves to determine the reversible work as the maximum work extracted from a system with the aid of a heat reservoir at temperature $T_{0}$ [8]. Moreover, the GB entropy is referred to as the same entropy as the one derived by Lieb and Yngvason in their axiomatic thermodynamics [8].
By taking the difference between two in Eq. (27), we get
$S_{2}-S_{1}=\frac{W_{0\to 2}^{\rm rev}-W_{0\to 1}^{\rm rev}-W_{1\to 2}^{\rm ad}}{T_{0}},$ (28)
@@
where $W_{1\to 2}^{\rm ad}=W_{1\to 0}^{\rm ad}+W_{0\to 2}^{\rm ad}$. Applying Kelvin principle to a cyclic process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\to A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$, we have
$W_{0\to 1}^{\rm rev}+W_{1\to 2}^{\rm ad}+W_{2\to 0}^{\rm rev}\leq 0\leftrightarrow-W_{0\to 1}^{\rm rev}-W_{1\to 2}^{\rm ad}+W_{0\to 2}^{\rm rev}\geq 0.$ (29)
@@
By using Eq. (29) to Eq. (28), we obtain the law of entropy non-decrease $S_{2}-S_{1}\geq 0$. Meanwhile, it may be considered that the entropy in Eq. (27) is dependent on the choice of the reference heat reservoir through the temperature $T_{0}$. In GB approach, it is shown that Eqs. (28) and (29) are in fact independent of $T_{0}$, and the entropy is a state quantity without calculating the reversible works in Eq. (27) explicitly, and any heat reservoir at any temperature can be used for the definition of entropy as the heat reservoir merely serves as an auxiliary system (see Sec. 7.4 of Ref. [5] for their proof). In contrast, in the present approach, we evaluated the reversible works explicitly by the use of the Carnot machine and showed that the entropy is independent of the choice of the reference heat reservoir (Eq. (10)).
@@
An advantage of GB entropy Eq. (27) to Clausius entropy Eq. (2) is that GB entropy is determined only from the measurement of $W_{0\to 1}^{\rm rev}$ for a reversible process with a heat reservoir at temperature $T_{0}$ and $W_{0\to 1}^{\rm ad}$ for an adiabatic process, with no need to measure heat. But, these two expressions are related through the expression Eq. (10).
3.2 Clausius entropy using isothermal heat
@@
By using the Clausius entropy Eq. (2) given $S_{0}$ as the entropy of a reference state, we can calculate the entropy of another state along any reversible path. In particular, when we choose a reversible path with a single heat reservoir at temperature $T_{R}=T^{\prime}$, we have the Clausius entropy in terms of an isothermal heat:
$S_{1}=S_{0}+\frac{Q_{{\rm iso},1}}{T^{\prime}},\ \ S_{2}=S_{0}+\frac{Q_{{\rm iso},2}}{T^{\prime}},$ (30)
@@
where $Q_{{\rm iso},1}\equiv\int_{0}^{1}\delta Q$ and $Q_{{\rm iso},2}\equiv\int_{0}^{2}\delta Q$. With this entropy in terms of the isothermal heat, it is possible to derive the law of entropy non-decrease Eq. (1) directly from the Kelvin principle by combining isothermal and adiabatic processes (see Sec. 8.11 of Ref. [9] for a similar argument): Consider that the equilibrium state $A_{1}=A_{1}(T_{1},V_{1})$ is brought to another equilibrium state $A_{2}=(T_{2},V_{2})$ adiabatically in an arbitrary manner. The equilibrium state $A_{2}$ is then brought to an equilibrium state $A_{2}^{\prime}=A_{2}^{\prime}(T^{\prime},V_{2}^{\prime})$ reversibly along the adiabatic curve. By attaching a prepared heat reservoir at temperature $T^{\prime}$ to the system, the equilibrium state $A_{2}^{\prime}$ is further brought to the equilibrium state $A_{1}^{\prime}=A_{1}^{\prime}(T^{\prime},V_{1}^{\prime})$ reversibly along the isothermal curve until it intersects the adiabatic curve with the original equilibrium state $A_{1}=A_{1}(T_{1},V_{1})$ on it. Finally, after detaching the heat reservoir from the system, the equilibrium state $A_{1}^{\prime}$ is brought to the original equilibrium state $A_{1}=A_{1}(T_{1},V_{1})$ reversibly along the adiabatic curve.
@@
The work extracted from the whole cycle process $A_{1}\to A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{2}^{\prime}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}^{\prime}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}$ is given as
$W_{\rm cyc}$ $=W_{1\to 2}^{\rm ad}+W_{2\to 2^{\prime}}^{\rm ad}+W_{2^{\prime}\to 1^{\prime}}^{\rm iso}+W_{1^{\prime}\to 1}^{\rm ad}$ (31)
$=U_{1}-U_{2}+U_{2}-U_{2}^{\prime}+W_{2^{\prime}\to 1^{\prime}}^{\rm iso}+U_{1}^{\prime}-U_{1}$
$=U_{1}^{\prime}-U_{2}^{\prime}+W_{2^{\prime}\to 1^{\prime}}^{\rm iso}$
$=U_{1}^{\prime}-U_{2}^{\prime}+U_{2}^{\prime}-U_{1}^{\prime}-T^{\prime}(S_{2}^{\prime}-S_{1}^{\prime})$
$=T^{\prime}(S_{1}^{\prime}-S_{2}^{\prime}),$
@@
where $W_{2^{\prime}\to 1^{\prime}}^{\rm iso}=U_{2}^{\prime}-U_{1}^{\prime}-T^{\prime}(S_{2}^{\prime}-S_{1}^{\prime})$ coincides with the Helmholtz free energy difference. By applying the Kelvin principle $W_{\rm cyc}=T^{\prime}(S_{1}^{\prime}-S_{2}^{\prime})=T^{\prime}(S_{1}-S_{2})\leq 0$, we derive the law of entropy non-decrease $S_{1}\leq S_{2}$.

### v5

@@ II.2 Derivation for simple systems
irrespective of whether the adiabatic process $A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{2}$ is reversible or irreversible.
Let $A_{0}$ be a reference equilibrium state of the system specified with $(T_{0},V_{0})$. We create the equilibrium state $A_{1}$ from the reference equilibrium state $A_{0}$ reversibly by combining two thermodynamic processes with the aid of a single heat reservoir at fixed temperature $T_{o}$ considered in Sec. II.1 (Fig. 2). The work extracted from this process is given by Eq. (11):
$W_{0\to 1}^{\rm rev}=U_{0}-U_{1}-T_{o}(S_{0}-S_{1}).$ (18)
@@
$W_{2\to 0}^{\rm rev}=-W_{0\to 2}^{\rm rev}=-(U_{0}-U_{2}-T_{o}(S_{0}-S_{2})).$ (19)
From Eqs. (17), (18), and (19), we obtain the work extracted from the whole cyclic process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$ as
$W_{\rm cyc}=W_{0\to 1}^{\rm rev}+W_{1\to 2}^{\rm ad}+W_{2\to 0}^{\rm rev}=T_{o}(S_{1}-S_{2}).$ (20)
By applying the Kelvin-Planck principle Eq. (5) to Eq. (20), we derive the law of entropy non-decrease Eq. (1).
II.3 Derivation for compound systems
We show Eq. (1) from Eq. (5) for a compound system. Here, a compound system is composed of $M$ subsystems in equilibrium, where each subsystem is separated by other subsystems by insulated walls. The state of $j$-th subsystem is specified with $(T_{j},V_{j})$ ($j=1,\cdots,M$). The internal energy, entropy, and volume of the compound system are given by the sum of those of subsystems as $U=\sum_{j=1}^{M}U_{j}$, and $S=\sum_{j=1}^{M}S_{j}$, and $V=\sum_{j=1}^{M}V_{j}$, respectively.
In the same manner as the simple system, we consider a cyclic process that consists of three thermodynamic processes $A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{2}$, $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}$, and $A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$ as follows. Let $A_{1}$ be the state of a compound system specified with the states of $M_{1}$ subsystems $\{(T_{1,1},V_{1,1}),(T_{1,2},V_{1,2}),\cdots,(T_{1,M_{1}},V_{1,M_{1}})\}$ in equilibrium, where $V_{1}=\sum_{j=1}^{M_{1}}V_{1,j}$ is the volume of the compound system at $A_{1}$. Consider an arbitrary adiabatic process $A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{2}$, where $A_{2}$ is specified with $M_{2}$ subsystems $\{(T_{2,1},V_{2,1}),(T_{2,2},V_{2,2}),\cdots,(T_{2,M_{2}},V_{2,M_{2}})\}$ in equilibrium. During this process, in addition to the adiabatic operations such as stirring, moving a piston attached to the system, and moving insulated walls partitioning subsystems, the number of subsystems is allowed to change from $M_{1}$ to $M_{2}$ by newly inserting or removing walls partitioning subsystems without any work cost. The work extracted from this process $W_{1\to 2}^{\rm ad}$ is given by
@@ II.3 Derivation for compound systems
Subsequently, we create the state $A_{1}$ from a reference equilibrium state $A_{0}$ of a simple system specified with $(T_{0},V_{0})$. First, we change the volume from $V_{0}$ to $V_{1}$ by a reversible adiabatic process, resulting in the temperature $T_{0^{\prime}}$, where we denote the state $(T_{0^{\prime}},V_{1})$ by $A_{0^{\prime}}$. The reversible work extracted from this process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0^{\prime}}$ is given by
@@
Then, we insert insulated walls (without any work cost) so that the system is divided into $M_{1}$ subsystems, where each subsystem is in the equilibrium state $A_{0^{\prime},j}$ specified with $(T_{0^{\prime}},V_{1,j})$ ($j=1,\dots,M_{1}$). By operating a Carnot machine between each subsystem and the heat reservoir at temperature $T_{o}$ until it reaches the equilibrium state with $(T_{1,j},V_{1,j})$, the total reversible work extracted from this process $A_{0^{\prime}}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}$ is given by
$W_{0^{\prime}\to 1}^{\rm rev}$ $=\sum_{j=1}^{M_{1}}(U_{0^{\prime},j}-U_{1,j}-T_{o}(S_{0^{\prime},j}-S_{1,j}))$ (23)
$=U_{0^{\prime}}-U_{1}-T_{o}(S_{0^{\prime}}-S_{1}).$
@@
Finally, the reversible work extracted from $A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$ should be calculated. Here, the system at the state $A_{2}$ consists of $M_{2}$ subsystems in equilibrium, whose state we denote by $A_{2,j}$ specified with $(T_{2,j},V_{2,j})$, where $\sum_{j=1}^{M_{2}}V_{j}=V_{2}$. In the similar manner to the calculation of $W_{0\to 1}^{\rm rev}$ in Eq. (24), we can calculate the reversible work extracted from $A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$ as the minus of $W_{0\to 2}^{\rm rev}$:
$W_{2\to 0}^{\rm rev}=-W_{0\to 2}^{\rm rev}=-(U_{0}-U_{2}-T_{o}(S_{0^{\prime\prime}}-S_{2})),$ (25)
@@
where $S_{0^{\prime\prime}}$ is the entropy at the state $A_{0^{\prime\prime}}$ specified with $(T_{0^{\prime\prime}},V_{2})$ reached from $A_{0}$ in a reversible adiabatic process. Note that $S_{0^{\prime}}=S_{0^{\prime\prime}}$ as $A_{0^{\prime}}$ and $A_{0^{\prime\prime}}$ (and $A_{0}$) are on the same adiabatic curve. By summing Eqs. (21), (24), and (25), and using $S_{0^{\prime}}=S_{0^{\prime\prime}}$, we obtain the work extracted from the whole cyclic process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$ as
$W_{\rm cyc}=W_{0\to 1}^{\rm rev}+W_{1\to 2}^{\rm ad}+W_{2\to 0}^{\rm rev}=T_{o}(S_{1}-S_{2}).$ (26)
@@
By applying the Kelvin-Planck principle in Eq. (5) to Eq. (26), we derive the law of entropy non-decrease in Eq. (1) for the compound system.
III Discussion
III.1 Gyftopoulos-Beretta entropy
@@
It should be noted that our formulation of entropy and its non-decrease in Sec. II may share some similarities with the approach by Gyftopoulos and Beretta (GB) [7] (see also Ref. [9]). By using notations in our formulation, their entropy is defined through the difference between works extracted from processes with and without the aid of a heat reservoir at fixed temperature $T_{o}$ [10]:
$S_{1}=S_{0}+\frac{W_{0\to 1}^{\rm rev}-W_{0\to 1}^{\rm ad/rev}}{T_{o}},\ S_{2}=S_{0}+\frac{W_{0\to 2}^{\rm rev}-W_{0\to 2}^{\rm ad/rev}}{T_{o}}.$ (27)
@@
where $W_{0\to 1}^{\rm ad/rev}$ and $W_{0\to 2}^{\rm ad/rev}$ are the works extracted from the reversible adiabatic processes. This definition is equivalent to Eq. (10) and is derived from it with the use of Eq. (6) by noting $W_{0\to 1}^{\rm ad/rev}=U_{0}-U_{1}$ for a reversible adiabatic process, but the one focusing on work rather than heat. This is a natural definition as Eq. (11) suggests that the entropy serves to determine the reversible work as the maximum work extracted from a system with the aid of a heat reservoir at temperature $T_{o}$ [8]. Moreover, it should be noted that the GB entropy is referred to as the same entropy as the one derived by Lieb and Yngvason in their axiomatic thermodynamics [8] (see also Ref. [11] for a related discussion).
By taking the difference between the two in Eq. (27), we get
$S_{2}-S_{1}=\frac{W_{0\to 2}^{\rm rev}-W_{0\to 1}^{\rm rev}-W_{1\to 2}^{\rm ad/rev}}{T_{o}},$ (28)
@@
where $W_{1\to 2}^{\rm ad/rev}=W_{1\to 0}^{\rm ad/rev}+W_{0\to 2}^{\rm ad/rev}$. Applying the Kelvin-Planck principle to a cyclic process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$, we have
$W_{0\to 1}^{\rm rev}+W_{1\to 2}^{\rm ad/rev}+W_{2\to 0}^{\rm rev}\leq 0\leftrightarrow-W_{0\to 1}^{\rm rev}-W_{1\to 2}^{\rm ad/rev}+W_{0\to 2}^{\rm rev}\geq 0.$ (29)
@@
By using Eq. (29) to Eq. (28), we obtain the law of entropy non-decrease $S_{2}-S_{1}\geq 0$. Meanwhile, it may be considered that the entropy in Eq. (27) is dependent on the choice of the reference heat reservoir through the temperature $T_{o}$. In GB approach, it is shown that Eqs. (28) and (29) are in fact independent of $T_{o}$: The GB entropy is shown to be a state quantity without an explicit calculation of the reversible works, and any heat reservoir at any temperature can be used for the definition of entropy as the heat reservoir merely serves as an auxiliary system (see Sec. 7.4 of Ref. [7] for their proof). In contrast, in the present approach, we evaluated the reversible works explicitly by the use of the Carnot machine and showed that the entropy is independent of the choice of the reference heat reservoir (Eq. (10)).
@@
An advantage of the GB entropy Eq. (27) to the Clausius entropy Eq. (2) is that the GB entropy is determined only from the measurement of $W_{0\to 1}^{\rm rev}$ for a reversible process with a heat reservoir at temperature $T_{o}$ and $W_{0\to 1}^{\rm ad/rev}$ for an adiabatic process, with no need to measure heat. But, these two expressions are related through the expression Eq. (10). See Table 1 for the summary of these three expressions.
Clausius (Eq. (2)) Gyftopoulos and Beretta [7] (Eq. (27)) This study (Eq. (10))
$S_{1}=S_{0}+\int_{0}^{1}\frac{\delta Q}{T_{R}}$ $S_{1}=S_{0}+\frac{W_{0\to 1}^{\rm rev}-(U_{0}-U_{1})}{T_{o}}$ $S_{1}=S_{0}+\int_{0}^{1}\frac{\delta Q}{T}$
Changeable temperature $T_{R}$, heat Fixed temperature $T_{o}$, work Fixed temperature $T_{o}$, heat
Table 1: Apparently different but equivalent three expressions of entropy: The Clausius expression uses heat divided by the changeable temperature of a heat reservoir; The Gyftopoulos-Beretta expression uses work divided by the fixed temperature of a single heat reservoir; The expression in this study uses heat divided by the system’s temperature while a single heat bath at fixed temperature is involved.
III.2 Clausius entropy using isothermal heat
@@
By using the Clausius entropy Eq. (2) given $S_{0}$ as the entropy of a reference state, we can calculate the entropy of another state along any reversible path. In particular, when we choose a reversible path that consists of isothermal processes with a single heat reservoir at temperature $T_{o}$ and adiabatic processes, we have the Clausius entropy in terms of isothermal heat:
$S_{1}=S_{0}+\frac{Q_{{\rm iso},1}}{T_{o}},\ \ S_{2}=S_{0}+\frac{Q_{{\rm iso},2}}{T_{o}},$ (30)
@@
where $Q_{{\rm iso},1}\equiv\int_{0}^{1}\delta Q$ and $Q_{{\rm iso},2}\equiv\int_{0}^{2}\delta Q$ for isothermal processes. With this entropy in terms of the isothermal heat, it is possible to derive the law of entropy non-decrease Eq. (1) directly from the Kelvin-Planck principle applied to the following cyclic process (see Sec. 8.11 of Ref. [12] for a similar argument): Consider that the equilibrium state $A_{1}=A_{1}(T_{1},V_{1})$ is brought to another equilibrium state $A_{2}=(T_{2},V_{2})$ adiabatically in an arbitrary manner. The equilibrium state $A_{2}$ is then brought to the equilibrium state $A_{2^{\prime}}=A_{2^{\prime}}(T_{o},V_{2^{\prime}})$ reversibly along the adiabatic curve. By attaching a prepared heat reservoir at temperature $T_{o}$ to the system, the equilibrium state $A_{2^{\prime}}$ is further brought to the equilibrium state $A_{1^{\prime}}=A_{1^{\prime}}(T_{o},V_{1^{\prime}})$ reversibly along the isothermal curve until it intersects the adiabatic curve with the original equilibrium state $A_{1}=A_{1}(T_{1},V_{1})$ on it. Finally, after detaching the heat reservoir from the system, the equilibrium state $A_{1^{\prime}}$ is brought to the original equilibrium state $A_{1}=A_{1}(T_{1},V_{1})$ reversibly along the adiabatic curve.
@@
The work extracted from the whole cyclic process $A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{2}\stackrel{{\scriptstyle\rm ad/rev}}{{\longrightarrow}}A_{2^{\prime}}\stackrel{{\scriptstyle\rm iso/rev}}{{\longrightarrow}}A_{1^{\prime}}\stackrel{{\scriptstyle\rm ad/rev}}{{\longrightarrow}}A_{1}$ is given as
$W_{\rm cyc}$ $=W_{1\to 2}^{\rm ad}+W_{2\to 2^{\prime}}^{\rm ad/rev}+W_{2^{\prime}\to 1^{\prime}}^{\rm iso/rev}+W_{1^{\prime}\to 1}^{\rm ad/rev}$ (31)
$=U_{1}-U_{2}+U_{2}-U_{2^{\prime}}+W_{2^{\prime}\to 1^{\prime}}^{\rm iso/rev}+U_{1^{\prime}}-U_{1}$
$=U_{1^{\prime}}-U_{2^{\prime}}+W_{2^{\prime}\to 1^{\prime}}^{\rm iso/rev}$
$=U_{1^{\prime}}-U_{2^{\prime}}+U_{2^{\prime}}-U_{1^{\prime}}-T_{o}(S_{2^{\prime}}-S_{1^{\prime}})$
$=T_{o}(S_{1^{\prime}}-S_{2^{\prime}}),$
@@
where $W_{2^{\prime}\to 1^{\prime}}^{\rm iso/rev}=U_{2^{\prime}}-U_{1^{\prime}}-T_{o}(S_{2^{\prime}}-S_{1^{\prime}})$ coincides with the minus of the Helmholtz free energy difference of the system before and after the isothermal process. By applying the Kelvin-Planck principle Eq. (5) to Eq. (31), we have $W_{\rm cyc}=T_{o}(S_{1^{\prime}}-S_{2^{\prime}})=T_{o}(S_{1}-S_{2})\leq 0$ and thus we derive the law of entropy non-decrease $S_{1}\leq S_{2}$.